OpenStudy (anonymous):

Please help! i will fan Choose the equation below that represents the line that passes through the point (2, 4) and has a slope of 3. y − 4 = 3(x − 2) y − 2 = 3(x − 4) y + 4 = 3(x + 2) y + 2 = 3(x + 4)

jimthompson5910 (jim_thompson5910):

Similar example: What is the equation of the line that has a slope of 9 and goes through the point (7,15) ? Slope = \(\Large \color{red}{9}\) The line goes through the point \(\Large (\color{green}{7},\color{blue}{15})\) \[\Large \color{red}{m} = \color{red}{9}\] \(\Large (\color{green}{x_1},\color{blue}{y_1})=(\color{green}{7},\color{blue}{15}) \rightarrow \color{green}{x_1}=\color{green}{7} \text{ and } \color{blue}{y_1}=\color{blue}{15}\) Point Slope form \[\Large y - y_1 = m(x - x_1)\] \[\Large y - \color{blue}{y_1} = \color{red}{m}(x - \color{green}{x_1})\] \[\Large y - \color{blue}{15} = \color{red}{9}(x - \color{green}{7})\]

jimthompson5910 (jim_thompson5910):

Another example: What is the equation of the line that has a slope of -2 and goes through the point (-1,5) ? Slope = \(\Large \color{red}{-2}\) The line goes through the point \(\Large (\color{green}{-1},\color{blue}{5})\) \[\Large \color{red}{m} = \color{red}{-2}\] \(\Large (\color{green}{x_1},\color{blue}{y_1})=(\color{green}{-1},\color{blue}{5}) \rightarrow \color{green}{x_1}=\color{green}{-1} \text{ and } \color{blue}{y_1}=\color{blue}{5}\) Point Slope form \[\Large y - y_1 = m(x - x_1)\] \[\Large y - \color{blue}{y_1} = \color{red}{m}(x - \color{green}{x_1})\] \[\Large y - \color{blue}{5} = \color{red}{-2}(x - (\color{green}{-1}))\] \[\Large y - 5 = -2(x + 1)\]